# Colors of Life in the Clouds: Biopigments of atmospheric microorganisms as a new signature to detect life on planets like Earth

### v2

@@ Abstract
When Carl Sagan and Ed Salpeter envisioned potential Sinkers, Floaters, and Hunters living in Jupiter’s clouds in 1976 (C. Sagan & E. E. Salpeter, 1976), the nature of life in Earth’s atmosphere remained widely unknown. Decades later, research has revealed a remarkable variety of microorganisms in our atmosphere. However, the spectral features of airborne microbes as biomarkers for detecting atmospheric life remained a mystery. Here, we present the first reflectance spectra of biopigments of atmospheric microorganisms based on laboratory cultivars of seven microbial strains isolated from Earth’s atmosphere. We show their distinct UV-resistant biosignatures and their impacts on models of diverse planetary scenarios, using Habitable Worlds Observatory (HWO) parameters. The reflectance of these biopigments from aerial bacteria creates the means to detect them on other Earth-like planets. It provides a paradigm shift that moves the search for life beyond the surface of a planet to ecosystems in atmospheres and clouds.
@@ 1.1 Microorganisms living in air
For some of these microorganisms, the atmosphere plays a crucial role in their life cycle, serving as a transport medium to disperse them over large distances. Part of the microorganisms may form an active microbial ecosystem (R. Lappan et al., 2024). Current technology cannot yet detect microbial life in our atmosphere using remote sensing equipment (A. C. Schuerger et al., 2018), mainly because of its low density.
@@
The survival of microorganisms in the atmosphere requires extreme tolerance to UV radiation, particularly above the ozone layer (35 km) (N. C. Bryan et al., 2019; T. Šantl Temkiv et al., 2022), desiccation (N. C. Bryan et al., 2019; T. Šantl Temkiv et al., 2022), oxidation stress (M. Vaïtilingom et al., 2013), and nutrient scarcity when outside of clouds. Organisms isolated from cloud aerosols show biopigments (N. C. Bryan et al., 2019; T. Šantl Temkiv et al., 2022); pigment production has been widely reported in biosurveys of the atmosphere (T. Šantl Temkiv et al., 2022, and references therein).
Biopigments are synthesized by various organisms, including bacteria, algae, and fungi, and are often crucial for surviving high incident sunlight exposure, such as that found in Earth’s atmosphere. Many fundamental questions about pigmented organisms and their place in the biosphere remain unanswered, given that colorful biodiversity patterns seem to relate more to the type of environment than to the organisms themselves (L. F. Coelho et al., 2022). Some pigments contribute remarkable resistance to UV irradiation, e.g., in polar and desert ecosystems (P. Vítek et al., 2017), where carotenoids produced by ice and snow biota play crucial roles in both photoprotection and light harvesting (L. F. Coelho et al., 2022).
Note that although diverse viable and stress-tolerant microorganisms have been reported in clouds on Earth, whether they are actively metabolizing while residing in the atmosphere remains an open question. But, even if the microorganism is no longer active, their biopigments can provide signatures of life on the surface (i.e., L. F. Coelho et al., 2022; S. Hegde et al., 2015; E. W. Schwieterman, 2018; L. F. Coelho et al., 2024), and in the atmosphere of exoplanets, especially for water-rich atmospheres.
@@
Here, we measure biopigments from organisms isolated from cloud aerosols and the atmosphere (e.g., N. C. Bryan et al., 2019), and show that UV-resistant pigments from microorganisms isolated from modern Earth atmospheres can change the search for life in clouds in other similar atmospheres.
1.2 Life in the atmosphere of exoplanets
To consider the range of evolutionary stages on exoplanets (e.g., J. T. O’Malley-James & L. Kaltenegger, 2018, 2019a; L. Kaltenegger et al., 2007, 2020), the search for extraterrestrial life focuses on two main approaches: i) searching for atmospheric chemical biosignature pairs (J. F. Kasting et al., 2014; L. Kaltenegger, 2017), such as the combination of oxygen and methane, and ii) searching for surface signatures, such as biopigments of microorganisms or vegetation in the spectrum (e.g., L. F. Coelho et al., 2022; S. Hegde et al., 2015; E. W. Schwieterman, 2018; L. F. Coelho et al., 2024; J. T. O’Malley-James & L. Kaltenegger, 2018, 2019a; S. Seager et al., 2005). Machine learning techniques can prioritize promising rocky exoplanet targets using initial identification of water and biological surfaces through color filters (e.g., D. Pham & L. Kaltenegger, 2021; S. Vannah et al., 2025), for upcoming telescopes like the Extremely Large Telescope (ELT) and concepts like HWO.
First detections of atmospheric signatures such as hydrogen, methane, oxygen, and carbon dioxide for exoplanets made by ground-based (e.g., F. Borsa et al., 2021) and space-based (e.g., JWST Transiting Exoplanet Community Early Release Science Team et al., 2023) telescopes, for some of the more than 6000 detected exoplanets, show diverse atmospheric compositions. Initial observations of rocky planets in the Habitable Zone have started but are not yet conclusive (e.g., O. Lim et al., 2023; E. Ducrot et al., 2024).
@@ 1.2 Life in the atmosphere of exoplanets
Until now, clouds have mostly been considered obstacles to identifying signs of life because they can obscure both their atmospheric and surface signatures (e.g., L. Kaltenegger et al., 2007), thus challenging both current atmosphere retrievals and future imaging. However, the possibility of a pigmented biosphere in the clouds of other Earth-like planets changes that narrative.
@@ 2.1 Reflectance measurements
We used an ASD FieldSpec 4 Spectrometer, which covers the wavelength range from 350–2500 nm at intervals of 1 nm, and an ASD integrating sphere to measure the reflectance of the samples. We deposited cultures homogeneously on a 25 mm plain white mixed cellulose ester filter (0.45 mm) using a 10 mL syringe and a filtration system to the point of filter saturation (10 mL ± 1 mL), corresponding to $10^{7}$ and $10^{8}$ cells per filter, as shown by S. Hegde et al. (2015). The maximum saturation approach provides an upper limit of signal strength as a baseline spectral reference for planetary modeling, while avoiding experimental artifacts of variable cell densities and ensuring reproducibility. While our spectra test whether such features are in principle detectable and represent maximal bioaerosol coverage, future work including these biopigments spectra into more complex atmosphere models, which could explore minimal cell concentrations for different aerosol distribution assumptions, atmospheric dynamics assumptions, and cloud coverage assumptions, could be used to assess the limit of detection for different instrument designs.
The same sample was measured at two different times: immediately after being deposited on the filter (fresh) and after 1 week of air drying in the dark at room temperature (dry) (L. F. Coelho et al., 2022). Filters with only culture media were measured as experimental controls, for both wet measurements (fresh medium in the filter) and dry measurements (dry medium in the filter). Reflectance spectra were presented in absolute values and values normalized to the average reflectance between 300 nm and 350 nm.
@@ 2.2 Data generation and simulations
We used the Exo-Prime II model (J. Madden & L. Kaltenegger, 2020) to generate high-resolution Earth-like atmospheric spectra with a minimum resolution of $\lambda/\Delta\lambda=100.000$. Ocean and snow surface albedo have been selected from the United States Geological Survey (USGS) Spectra Library (R. N. Clark & others, 2003), and the single-layer cloud albedo is based on the 20 mm MODIS model (M. King et al., 1997; W. B. Rossow & R. A. Schiffer, 1999) following previous work (J. Madden & L. Kaltenegger, 2020). We resampled the high-resolution spectra of the atmosphere reflectance of an Earth-like atmosphere and the lower-resolution biota (Figure 2), cloud, ocean, and snow albedos using the Python Spectres 2.2.0 package (A. C. Carnall, 2017). We resampled the low-resolution biota, cloud, ocean, and snow albedos to R = 140 (comparable to the HWO concept). Our exoplanet models contain reflectance spectra of frozen worlds and snowball planets, Earth-analogues, and ocean worlds with a different surface coverage of purple bacteria for atmospheres with and without clouds.
@@ 3 Results and Discussion
Table 1: Absorption features (nm) from fresh and dry samples. Taxonomy is represented by 16S rRNA gene Sequence Identity.
@@
Figure 1: Relative reflectance spectra of seven aerial microorganisms of wet (top frame) and dry (middle frame) measurements, and absolute reflectance (dashed lines for dry biota; solid lines for wet biota) with zoomed-in visible range wavelength panel on the right. Control (black lines in the bottom frame) refers to a filter with culture medium only, where wet control refers to fresh medium and dry control refers to dry medium. Biopigments show features between 400 and 600 nm (see Zoom), water absorption features at 1490 nm and 1900 nm are indicated by vertical blue bands. (Dataset available on Zenodo).
@@
Looking closely at the individual signatures from the different samples, we see distinctive spectral fingerprints from these microorganisms. L9-9A-1 reflectance signatures at 410 nm, 480 nm, 505 nm, and 540 nm agree with the absorption pattern of spirilloxanthin (E. Papagiannakis et al., 2003), a carotenoid known for transferring the photon energy absorbed at 540 nm and dissipating this excess energy as heat (R. G. Saer & R. E. Blankenship, 2017). L9-4 and L6-1 spectra show the presence of pink/red/orange carotenoids such as beta-carotene or lycopene (absorption features at 420 nm, 470 nm, 490 nm; F. A. D. Sousa et al., 2014). The yellow pigments from L7-1, L7-7A, L7-5, and L9-7 match with patterns of carotenoids zeaxanthin and lutein described in the literature with absorption features at 410 nm, 440 nm, 430 nm, 450 nm, 460 nm, and 480 nm (C. N. N’soukpoé-Kossi & R. M. Leblanc, 1988), as well as flavins (absorption features at 450 nm; P. F. Heelis, 1982). Flavins are electron-carrying yellow pigments produced by microorganisms related to several metabolic pathways, usually linked to growth strategies. Note that small wavelength shifts in the literature are likely solvent-dependent. We are measuring the whole cell and not the biopigment dissolved in an organic solvent.
@@
Dry biopigments in Figure 1 (bottom frame) show an increase in reflectivity (L. F. Coelho et al., 2022), allowing a better observation of the yellow and fainter pigments from L7-1, L7-7A, and L7-5. Dry measurements are analogues of drier environments than a typical humid ecosystem, such as ice or non-polar deserts. In these environments, biopigments will reflect strongly (higher intensity), and the increment will be proportional to the water loss (S. Hegde et al., 2015; L. F. Coelho et al., 2024).
@@
The features of the observed carotenoids (absorption features and reflectance peaks – Table 1) are more distinct in the wet measurements (Figure 1, top frame). This result suggests that atmospheric signatures of carotenoids are easier to distinguish in aerial environments where water is more available because the signature features are more pronounced. However, the overall intensity of the reflectivity of the microorganisms could be higher in drier atmospheres.
@@
Figure 2: Modeled reflectance spectra of different Earth-like planets: Ocean Worlds (left), Snowball Planets (right), both with clouds covered by strain L9-9A-1 (see Table 1 and Figure 1), resampled for an HWO resolution of R=140. Purple lines represent planets with aerial biota in clouds. Top row: 100% clouds with 100% covered by aerial biota. Middle row: 100% clouds with 50% covered by aerial biota. Bottom row: 50% clouds with 50% covered by aerial biota. Dashed grey lines show “lifeless cloudy planet”, solid blue shows “lifeless ocean planet”, and solid pink shows “lifeless snowball planet” scenarios. Individual spectra of albedos in Figure 3.
@@
To explore the effect of the biota on the observable reflection spectrum of exoplanets (Figure 1), we show the six model scenarios of planets with different surfaces, with resolution resampled to R = 140, comparable to the design concepts of HWO in the early LUVOIR report (Figure 2). We simulate planetary surfaces covered with oceans or snow, a modern Earth-like atmosphere, and varying cloud coverage (100% and 50%) with different fractions of clouds covered with biota (100% and 50%). Note that especially for fully cloud-covered planets in harsh UV environments like around M-stars (e.g., J. Scalo et al., 2007; J. T. O’Malley-James & L. Kaltenegger, 2019b), producing pigments for UV protection should be evolutionarily favorable and increase the detectability of life in the atmosphere. While Earth currently is not in such an environment in the Solar System, it is unknown whether a cloud-covered planet could show a full or 50% coverage of biota, which could be detected observationally as shown by the model scenarios (Figure 2). However, while full cloud coverage on an exoplanet generally decreases the possibility of finding surface life through reflection measurements, life in the atmosphere could leave a strong signal, shifting the search for life to the pigments that can be detected in the atmosphere, indicating surface life that is hidden from view.
@@
Adding clouds generally increases the reflectivity and detectability of the exoplanet in the visible compared to an ocean-covered surface because of their higher reflectivity, but decreases the overall reflectivity for snowball planets because of the lower reflectivity of clouds compared to snow (see individual albedos in Figure 3). Once biota is added to the clouds (see Figure 2), it modifies the shape of the reflection spectrum, especially around 500 nm, where the pigment absorption of aerial biota shows strong features for both wet and dry biota (see Figure 1). Overall, wet biota reduces the reflectivity of clouds (Figure 2 left, ocean worlds) and dry biota increases the reflectivity of clouds (Figure 2 right, snowball planets). The interplay between biotic and abiotic components in cloud layers suggests that inhabited clouds would leave a detectable imprint on the planet’s overall spectral signature, providing a potential biosignature for future exoplanet observations with missions like HWO.

### v3

@@ Abstract
When Carl Sagan and Ed Salpeter envisioned potential sinkers, floaters, and hunters living in Jupiter’s clouds in 1976 (C. Sagan & E. E. Salpeter, 1976), the nature of life in Earth’s atmosphere remained widely unknown. Decades later, research has revealed a remarkable variety of microorganisms in our atmosphere. However, the spectral features of airborne microbes as biomarkers for detecting atmospheric life remained a mystery. Here, we present the first reflectance spectra of biopigments of atmospheric microorganisms based on laboratory cultivars of seven microbial strains isolated from Earth’s atmosphere. We show their distinct UV-resistant biosignatures and their impacts on models of diverse planetary scenarios, using Habitable Worlds Observatory parameters. The reflectance of these biopigments from aerial bacteria creates the means to detect them on other Earth-like planets. It provides a paradigm shift that moves the search for life beyond the surface of a planet to ecosystems in atmospheres and clouds.
@@ 1.1 Microorganisms living in air
For some of these microorganisms, the atmosphere plays a crucial role in their life cycle, serving as a transport medium to disperse them over large distances. Part of the microorganisms may form an active microbial ecosystem (R. Lappan et al., 2024). Current technology cannot yet detect microbial life in our atmosphere using remote-sensing equipment (A. C. Schuerger et al., 2018), mainly because of its low density.
@@
The survival of microorganisms in the atmosphere requires tolerance to UV radiation, particularly above the ozone layer at 35 km (N. C. Bryan et al., 2019; T. Šantl Temkiv et al., 2022), desiccation (N. C. Bryan et al., 2019; T. Šantl Temkiv et al., 2022), oxidation stress (M. Vaïtilingom et al., 2013), and nutrient scarcity when outside of clouds. Organisms isolated from cloud aerosols show biopigments (N. C. Bryan et al., 2019; T. Šantl Temkiv et al., 2022); pigment production has been widely reported in biosurveys of the atmosphere (T. Šantl Temkiv et al., 2022, and references therein).
Biopigments are synthesized by various organisms, including bacteria, algae, and fungi, and are often crucial for surviving high-incident sunlight exposure, such as that found in Earth’s atmosphere. Many fundamental questions about pigmented organisms and their place in the biosphere remain unanswered, given that colorful biodiversity patterns seem to relate more to the type of environment than to the organisms themselves (L. F. Coelho et al., 2022). Some pigments contribute remarkable resistance to UV irradiation, e.g., in polar and desert ecosystems (P. Vítek et al., 2017), where carotenoids produced by ice and snow biota play crucial roles in both photoprotection and light harvesting (L. F. Coelho et al., 2022).
Note that although diverse viable and stress-tolerant microorganisms have been reported in clouds on Earth, whether they are actively metabolizing while residing in the atmosphere remains an open question. But, even if the microorganism is no longer active, their biopigments can provide signatures of life on the surface (e.g., L. F. Coelho et al., 2022; S. Hegde et al., 2015; E. W. Schwieterman, 2018; L. F. Coelho et al., 2024), and in the atmosphere of exoplanets, especially for water-rich atmospheres.
@@
Here, we measure biopigments from organisms isolated from cloud aerosols and the atmosphere (e.g., N. C. Bryan et al., 2019), and show that UV-resistant pigments from microorganisms isolated from Earth’s atmosphere can refocus the search for life toward the clouds on Earth-like planets.
1.2 Life in the atmospheres of exoplanets
To consider the range of evolutionary stages on exoplanets (e.g., J. T. O’Malley-James & L. Kaltenegger, 2018, 2019a; L. Kaltenegger et al., 2007, 2020), the search for extraterrestrial life focuses on two main approaches: i) searching for atmospheric chemical biosignature pairs (J. F. Kasting et al., 2014; L. Kaltenegger, 2017), such as the combination of oxygen and methane, and ii) searching for surface signatures, such as biopigments of microorganisms or vegetation in the spectrum (e.g., L. F. Coelho et al., 2022; S. Hegde et al., 2015; E. W. Schwieterman, 2018; L. F. Coelho et al., 2024; J. T. O’Malley-James & L. Kaltenegger, 2018, 2019a; S. Seager et al., 2005). Machine learning techniques can prioritize promising rocky exoplanet targets using initial identification of water and biological surfaces through color filters (e.g., D. Pham & L. Kaltenegger, 2021; S. Vannah et al., 2025), for upcoming telescopes like the Extremely Large Telescope and concepts like the Habitable Worlds Observatory (HWO).
First detections of atmospheric signatures such as hydrogen, oxygen, methane, and carbon dioxide for exoplanets made by ground-based (e.g., A. G. Jensen et al., 2012; F. Borsa et al., 2021) and space-based (e.g., T. J. Bell et al., 2023; JWST Transiting Exoplanet Community Early Release Science Team et al., 2023) telescopes, for some of the more than 6000 detected exoplanets, show diverse atmospheric compositions. Initial observations of rocky planets in the habitable zone have started but are not yet conclusive (e.g., O. Lim et al., 2023; E. Ducrot et al., 2024).
@@ 1.2 Life in the atmospheres of exoplanets
Until now, clouds have mostly been considered obstacles to identifying signs of life because they can obscure both their atmospheric and surface signatures (e.g., L. Kaltenegger et al., 2007; S. Kelkar et al., 2025), thus challenging both current atmosphere retrievals and future imaging. However, the possibility of a pigmented biosphere in the clouds of other Earth-like planets changes that narrative.
@@ 2.1 Reflectance measurements
We used an ASD FieldSpec 4 Spectrometer, which covers the wavelength range from 350–2500 nm at intervals of 1 nm, and an ASD integrating sphere to measure the reflectance of the samples. We deposited cultures homogeneously on a 25 mm plain white mixed cellulose ester filter (0.45 mm) using a 10 mL syringe and a filtration system to the point of filter saturation (10 mL ± 1 mL), corresponding to $10^{7}$ and $10^{8}$ cells per filter, as shown by S. Hegde et al. (2015). The maximum saturation approach provides an upper limit of signal strength as a baseline spectral reference for planetary modeling while avoiding experimental artifacts of variable cell densities and ensuring reproducibility.
The same sample was measured at two different times: immediately after being deposited on the filter (wet) and after 1 week of air drying in the dark at room temperature (dry) (L. F. Coelho et al., 2022). Filters with only culture media were measured as experimental controls, for both wet measurements (fresh medium in the filter) and dry measurements (dry medium in the filter). Reflectance spectra were presented in absolute values and values normalized to the average reflectance between 300 nm and 350 nm.
@@ 2.2 Data generation and simulations
We used the Exo-Prime II model (J. Madden & L. Kaltenegger, 2020) to generate high-resolution Earth-like atmospheric spectra with a minimum resolution of $\lambda/\Delta\lambda=100.000$. Ocean and snow surface albedos were selected from the United States Geological Survey Spectra Library (R. N. Clark & others, 2003), and the single-layer cloud albedo is based on the 20 mm MODIS model (M. King et al., 1997; W. B. Rossow & R. A. Schiffer, 1999) following previous work (J. Madden & L. Kaltenegger, 2020). We resampled the high-resolution spectra of the atmosphere reflectance of an Earth-like atmosphere and the lower-resolution biota (Figure 2), cloud, ocean, and snow albedos using the Python Spectres 2.2.0 package (A. C. Carnall, 2017). We resampled the low-resolution biota, cloud, ocean, and snow albedos to R = 140 (comparable to the HWO concept). Our exoplanet models contain reflectance spectra of snowball planets, and ocean worlds with a different surface coverage of cloud biota for atmospheres with 100% and 50% clouds.
@@ 3 Results and Discussion
Table 1: Absorption features (nm) of biopigmetns from wet and dry biota samples. Taxonomy is represented by 16S rRNA gene Sequence Identity.
@@
Figure 1: Relative reflectance spectra of seven aerial microorganisms of wet (top panel) and dry (middle panel) measurements, and absolute reflectance (bottom panel - dashed lines for dry biota; solid lines for wet biota) with zoomed-in visible range wavelength panel on the right. Control (black lines in the bottom panel) refers to a filter with culture medium only, where wet control refers to fresh medium and dry control refers to dry medium. Biopigments show features between 400 and 600 nm (see ”Zoom” panels), water absorption features at 1490 nm and 1900 nm are indicated by vertical blue bands. (Data set available on Zenodo).
@@
Looking closely at the individual signatures from the different samples, we see distinctive spectral fingerprints from these microorganisms. L9-9A-1 reflectance signatures at 410, 480, 505, and 540 nm agree with the absorption pattern of spirilloxanthin (E. Papagiannakis et al., 2003), a carotenoid known for transferring the photon energy absorbed at 540 nm and dissipating this excess energy as heat (R. G. Saer & R. E. Blankenship, 2017). L9-4 and L6-1 spectra show the presence of pink/red/orange carotenoids such as beta-carotene or lycopene (absorption features at 420, 470, 490 nm; F. A. D. Sousa et al., 2014). The yellow pigments from L7-1, L7-7A, L7-5, and L9-7 match with patterns of carotenoids zeaxanthin and lutein described in the literature with absorption features at 410 nm, 440 nm, 430, 450, 460, and 480 nm (C. N. N’soukpoé-Kossi & R. M. Leblanc, 1988), as well as flavins (absorption features at 450 nm; P. F. Heelis, 1982). Flavins are electron-carrying yellow pigments produced by microorganisms related to several metabolic pathways, usually linked to growth strategies. Note that small wavelength shifts in the literature are likely solvent dependent. We are measuring the whole cell and not the biopigment dissolved in an organic solvent.
@@
Dry biopigments in Figure 1 (bottom panel) show an increase in reflectivity (L. F. Coelho et al., 2022), allowing a better observation of the yellow and fainter pigments from L7-1, L7-7A, and L7-5. Dry measurements are analogs of drier environments than a typical humid ecosystem, such as ice or nonpolar deserts. In these environments, biopigments will reflect strongly (higher intensity), and the increment will be proportional to the water loss (S. Hegde et al., 2015; L. F. Coelho et al., 2024).
@@
The features of the observed carotenoids (absorption features and reflectance peaks – Table 1) are more distinct in the wet measurements (Figure 1, top panel). This result suggests that atmospheric signatures of carotenoids are easier to distinguish in aerial environments where water is more available because the signature features are more pronounced. However, the overall intensity of the reflectivity of the microorganisms could be higher in drier atmospheres.
@@
Figure 2: Modeled reflectance spectra of different Earth-like planets: ccean worlds (left column), snowball planets (right column), both with clouds covered by strain L9-9A-1 (see Table 1 and Figure 1), resampled for an HWO resolution of R=140. Purple lines represent planets with aerial biota in clouds. Top row: 100% clouds with 100% covered by aerial biota. Middle row: 100% clouds with 50% covered by aerial biota. Bottom row: 50% clouds with 50% covered by aerial biota. Dashed gray lines show “lifeless cloudy planet”, solid blue shows “lifeless ocean planet”, and solid pink shows “lifeless snowball planet” scenarios. Atmospheric features are labeled O2 at 690, 760, and 1260 nm, and H2O at 1490 and 1900 nm. Individual spectra of albedos are shown in Figure 3.
@@
To explore the effect of the biota on the observable reflection spectrum of exoplanets (Figure 1), we show the six model scenarios of planets with different surfaces, with resolution resampled to R = 140, comparable to the design concepts of HWO in the early LUVOIR report (Figure 2). We simulate planetary surfaces covered with oceans or snow, a modern Earth-like atmosphere, and varying cloud coverage (100% and 50%) with different fractions of clouds covered with biota (100% and 50%). Note that especially for fully cloud-covered planets in harsh UV environments like around M stars (e.g., J. Scalo et al., 2007; J. T. O’Malley-James & L. Kaltenegger, 2019b), producing pigments for UV protection should be evolutionarily favorable and increase the detectability of life in the atmosphere. While Earth currently is not in such an environment in the solar system, it is unknown whether a cloud-covered planet could show a full or 50% coverage of biota, which could be detected observationally as shown by the model scenarios (Figure 2). However, while full cloud coverage on an exoplanet generally decreases the possibility of finding surface life through reflection measurements, life in the atmosphere could leave a strong signal, shifting the search for life to the pigments that can be detected in the atmosphere, indicating surface life that is hidden from view.
@@
Adding clouds generally increases the reflectivity and detectability of the exoplanet in the visible compared to an ocean-covered surface because of their higher reflectivity but decreases the overall reflectivity for snowball planets because of the lower reflectivity of clouds compared to snow (see individual albedos in Figure 3). Once biota is added to the clouds (see Figure 2), it modifies the shape of the reflection spectrum, especially around 500 nm, where the pigment absorption of aerial biota shows strong features for both wet and dry biota (see Figure 1). Overall, wet biota reduces the reflectivity of clouds (Figure 2 left, ocean worlds), and dry biota increases the reflectivity of clouds (Figure 2 right, snowball planets). The interplay between biotic and abiotic components in cloud layers suggests that inhabited clouds would leave a detectable imprint on the planet’s overall spectral signature, providing a potential biosignature for future exoplanet observations with missions like HWO.
While our work represents maximal bioaerosol coverage, future work, including this reference spectra into more complex atmospheric models, could explore minimal cell concentrations for different aerosol distribution assumptions, atmospheric dynamics assumptions, and cloud coverage assumptions. This will be important to assess the sensitivity and limit of detection for different instrument designs.